(a) Find the inverse of 19 modulo 141, using the Extended Euclidean Algorithm.


Show your steps.


(b) Solve the congruence 19x ≡ 7 (mod 141), by specifying all the integer solutions x


that satisfy the congruence.


1
Expert's answer
2021-11-02T18:25:39-0400

(a) Let us find the inverse of 19 modulo 141, using the Extended Euclidean Algorithm. Since

"141=19\cdot 7+8,\\19=8\cdot 2+3,\\ 8=3\cdot 2+2,\\ 3=2\cdot 1+1,\\ 2=1\cdot 2+0,"

we get that

"1=3-2\cdot 1= 3-(8-3\cdot 2)=-8+3\cdot 3\\=-8+3(19-8\cdot 2)
=3\cdot 19-8\cdot 7\\=
19\cdot 3-(141-19\cdot 7)7=141(-7)+19\cdot 52."

Therefore, "141(-7)+19\cdot 52\equiv1(\mod 141)," and hence

"19\cdot 52\equiv1(\mod 141)."

We conclude that 52 is the inverse of 19 modulo 141.


(b) Let us solve the congruence 19x ≡ 7 (mod 141),

Taking into account that "19\cdot 52\equiv1(\mod 141)," we conclude that "19\cdot (52\cdot 7)\equiv 7(\mod 141)," and thus

"x\equiv 52\cdot 7(\mod 141)\equiv 364(\mod 141)\equiv 82(\mod 141)."

We conclude that the solution is of the form

"x=82+141k," where "k\in\Z."